# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v1

@@ 4 Theoretical analysis
BBoxER, like all post-training approaches, raises foundational concerns regarding its ability to preserve generalization, its vulnerability to poisoning, its potential to leak private information, and its susceptibility to misuse for bypassing access restrictions. We address these aspects within a unified framework, analyzing retrofitted models through the lenses of generalization, privacy, and robustness. For mathematical background see App. C, additional results in App. E and proofs in App. G.
Formally, let $S(\omega,D,a,b)$ denote the internal state $I_{b}$ of a learning algorithm $a$, after $b$ iterations on dataset $D$ with randomness $\omega$. The number of possible internal states over all datasets is $N(\omega,a,b)=\operatorname{Card}\{S(\omega,D,a,b)\mid D\in{\cal D}\}$, which can vary even for fixed $\omega$ due to the variability of $D$. Let $k_{i}(\omega,a,b)$ be the number of possible values extracted from $D$ at iteration $i$; then, as shown in 5 of Alg. 1, we have
@@
4.1 Overfitting bounds
Retrofitting on a small budget raises the question of whether the updated model still generalizes or merely overfits the few examples used for post-training. We address this by bounding the generalization gap – that is, the difference between empirical and true loss – in terms of the variability of the algorithm’s internal state. Notably, these bounds are governed by the optimization path, with no dependence on the number of model parameters optimized during the retrofitting process.
Let $L$ denote the generalization error for a dataset distribution $F\in{\cal F}$ and $\widehat{L}$ the empirical error on $D$. We write $\widehat{L}(x)$ and $L(x)$ as shorthand for $\widehat{L}(modified(m_{0},x))$ and $L(modified(m_{0},x))$. No i.i.d. assumption is needed for $D$, only concentration inequalities. We fix the optimization randomness $\omega$ (the stochastic case is discussed in Sec. E.2) and study how the number of distinct possible internal states affects generalization. In Thm 1, deferred to App. E, we show that the generalization gap of the final model $\widehat{x}$ is controlled by the number of reachable configurations in  Equation 4. It also provides a uniform bound over all intermediate models $(m_{1},\dots,m_{b})$, ensuring that empirical losses remain close to the true risk throughout optimization.
The assumptions for  Thm 1 are mild: it only requires that empirical loss (on one draw of the dataset $D$) is a good proxy for population (test) performance, a standard concentration-bound hypothesis in classical learning theory [77, 78, 79, 61] (see  Sec. C.1). When losses are bounded in $[0,1]$ (as is the case in our experiments, where we use average accuracy),  Thm 1 with Hoeffding’s inequality yields:
@@ 4.1 Overfitting bounds
(informal) Under standard concentration assumptions (Equation A.1), if individual losses are in $[0,1]$ and Hoeffding’s inequality applies (Eq. A.2), then:
$P_{\omega_{D}}\left(|\widehat{L}(\widehat{x})-L(\widehat{x})|\geq\epsilon\right)\leq 2\cdot(\Pi_{i}k_{i})\cdot\exp(-2s\epsilon^{2})=2\cdot\exp\left(\ln 2\cdot\sum_{i=1}^{b}\log_{2}(k_{i})-2s\epsilon^{2}\right)$ (5)
As discussed in Secs. 3.2 and D, most BBO algorithms used within BBoxER bound the average $k_{i}$ by $2$ or $3$. Interpreting $\log_{2}(k_{i})$ as a budget in bits, this means that to avoid overfitting, we can afford a number of iterations $b$ proportional to $s$ divided by the bits used per step.
Bounds on additional BBO algorithms. Our analysis extends to a wide range of evolutionary BBO methods. In particular, Sec. E.4 shows that Bet-And-Run can reduce overfitting by distributing the budget over $N$ parallel runs, effectively scaling the total budget while preserving generalization. Sec. E.5 and Sec. E.6 provide analogous bounds for Population-Based Evolution Strategies (e.g., CMA, DiagonalCMA), Differential Evolution, and Particle Swarm Optimization.
@@ 4.2 Privacy and poisoning
Retrofitting modifies a model post-hoc based on user feedback, such as satisfaction or preference scores. While this enables iterative improvement, it also introduces vulnerabilities. Two key risks are privacy leakage, where individual feedback might be inferred from the final model, and poisoning, where a few corrupted inputs can steer the outcome. Both concerns relate to a notion of robustness to localized perturbations in user feedback. We first consider a simplified scenario: a binary comparison between two models based on majority vote over $n$ independent user preferences $r_{i}\in\{0,1\}$, with $r_{i}=1$ indicating preference for the first model. The observed frequency $f=\frac{1}{n}\sum_{i=1}^{n}r_{i}$ estimates an unknown true preference probability $f_{0}$.
@@
An attacker may corrupt up to $k$ values among the $r_{i}$, resulting in a perturbed frequency $f^{\prime}$ satisfying $|f^{\prime}-f|\leq k/n$. This setup follows the standard poisoning model [80], where the adversary is computationally unbounded and fully informed, but constrained in the number of values it can alter. Lem. 1 in App. G, bounds the probability that such a perturbation changes the comparison outcome.
@@
We now extend this to the full retrofitting algorithm (Alg. 1), which performs $b$ such comparisons. At iteration $j\in\{1,\dots,b\}$, the algorithm collects binary preferences $r_{j,1},\dots,r_{j,n}$, aggregates them into a frequency $f_{j}=\frac{1}{n}\sum_{i=1}^{n}r_{j,i}$, and selects a model based on whether $f_{j}>1/2$. Under poisoning, each $f_{j}$ may be shifted to $f^{\prime}_{j}$, with $|f^{\prime}_{j}-f_{j}|\leq k/n$, and we assume the adversary perturbs each round independently. The following theorem bounds the impact of these perturbations on the final output:
Theorem 2 (Privacy and robustness to poisoning).
Consider $output_{1}$ the resulting model $\texttt{FinalModel}=\texttt{modified}(m_{0},\hat{x})$ based on decisions $choice_{j}=\text{sign}(f_{j}-\tfrac{1}{2})$, and $output_{2}$ the output using perturbed frequencies $f^{\prime}_{j}$ satisfying $|f^{\prime}_{j}-f_{j}|\leq k/n$. Then:
$P\left(output_{1}\neq output_{2}\right)\leq b\frac{2k+1}{\sqrt{2n\pi}}.$ (6)
This bound relies on a union bound argument and the same adversarial model as [80]. Given two datasets $D_{1}$ and $D_{2}$ differing by one prompt or answer or any data unrelated to preference signals, then the outcome of Alg. 1 remains unchanged. As a result, the algorithm behaves identically on both datasets, satisfying perfect differential privacy (i.e. $\epsilon=0,\text{ and }\delta=0$ in Equation A.11 in Sec. F.1). This result has then tremendous consequences for poisoning and privacy. It shows that, unless $k$ is of order $\sqrt{n}$, the retrofitting process is unlikely to be affected by poisoning. And it ensures privacy by design, by guaranteeing the privacy of queries and responses as well as that of the user satisfaction signal. These implications are discussed in more detail in  App. F.
4.3 Robustness to extraction attacks from compression
@@
Formally, let $m_{0}$ be fine-tuned on unrelated datasets $D_{1},\dots,D_{n}$. For text, "unrelated" can mean differing in roughly one word out of five while remaining coherent. A dataset $D$ with $B$ bits may yield $n\approx 2^{B/k}$ such variants by altering one bit per $k$. We define vulnerability to data extraction as follows:
@@
This criterion can be tested practically: generate $n$ unrelated variants $D_{1},\dots,D_{n}$, fine-tune models on each $D_{i}$, and check whether each model assigns the highest likelihood to its own training set. If so, it indicates vulnerability to data extraction. We argue this is unlikely for retrofitting, due to limited capacity to encode training data. Thm 5 in Sec. G.2 formalizes this, leading to the corollary below:
@@
For given $\omega,a,b$ s.t. $k_{i}\leq 2$ for all $i$, BBoxER is not vulnerable to extraction for ${\cal D}=(D_{1},\dots,D_{n})$ if $n>2^{b}$. More generally, BBoxER is not vulnerable for ${\cal D}=(D_{1},\dots,D_{n})$ if $n>\prod_{i}k_{i}$.
Thus, if the number of alternative datasets exceeds that of distinguishable outputs, no extractor can identify the training data. Since $2^{b}$ grows only exponentially with budget $b$, while the number of unrelated textual variants grows exponentially with dataset size (e.g., $>400$B bits for Wikipedia articles, $>20$M for War and Peace), extraction is unlikely in practice.
4.4 Practical implications
Our bounds apply to a broad class of algorithms. As shown in Tab. 3 (Sec. D.1), most BBO methods have low average branching factors, typically bounded by two, ensuring strong guarantees on robustness, privacy, and overfitting.
Robustness to extraction attacks holds when the budget $b$ is much smaller than the dataset size (Sec. 4.3); robustness to poisoning is preserved up to $k\ll\sqrt{n}/b$ corrupted preferences, where $n$ is the total number of comparisons (Sec. 4.2); privacy remains high in all regimes; and overfitting is controlled for budgets $b$ linear in dataset size $s$ (Sec. 4.1), with bounds independent of model size. Bet-And-Run strategies further reduce overfitting via diversification (Sec. E.4), effectively increasing usable budget by the number of parallel runs. Sec. E.7 gives explicit thresholds under which generalization bounds remain meaningful.
@@ 5 Experimental results: Learning with Privacy and Robustness
The theoretical bounds in Sec. 4 establish that comparison based BBOs in the framework of BBoxER offer strong generalization guarantees and inherent privacy. However, theoretical guarantees alone do not ensure that these methods are practically effective—only empirical evidence can reveal whether they actually lead to meaningful learning. To validate the capacity of BBoxER as an LLM retrofitting method, we focus on verifiable domains, like math, where performance improvements are both measurable and interpretable. We can perform experiments with budgets $b$ determined theoretically for guaranteed generalization to the test set. This setup also allows us to empirically verify how performance changes when the budget increases beyond these thresholds, and how it transfers out of domain, from one task to another.
@@ 5.1 Experiments
Experimental setup. We deploy a family of BBO algorithms with finite branching factors, fully described in App. D using the Nevergrad implementation [46]. We conduct experiments on the Llama 3.1–8B [57] and Qwen 2.5–3B [58] models, applying BBoxER to both base and instruct variants. Retrofitting the base models allows us to verify that learning indeed occurs, while targeting the instruct models demonstrates that our method can yield further improvements even on top of models already fine-tuned for reasoning tasks. While we run multiple seeds where feasible, some experiments rely on a single seed due to the high computational cost—consistent with common practice in LLM research. All experimental details can be found in App. H.
Datasets. Most of our experiments deploy BBoxER on the GSM8K  [52], a dataset composed of short grade-school math problems, consisting of $7,473$ problems for training and $1,319$ for testing. We train on the train split and test on GSM8k-test (in distribution) with standard 8-shot methodology (unless otherwise stated), in addition to various other benchmark datasets (out of distribution): MATH500 (4-shots)  [53], Hellaswag (0-shots)  [54], GSM+ test_mini (8-shots)  [55], ARC Easy (0-shots) and ARC Challenge (0-shots)  [56], AMC23 (0-shots). All benchmarks are evaluated with exact match using greedy decoding and a single generation. The theoretical bounds in Sec. 4 allow us to precisely determine the budget $b$ that guarantees in-domain generalization (as a function of dataset size): for GSM8k-train the permissible budget is between $100$-$150$ steps for ”pure” BBO and $\approx 3$ times larger for Bet-and-run variants like 3-OneFifth. Some experiments probe beyond this budget.
Choices for parameter updates. A key design choice in our framework is determining which parameters the BBO algorithm should probe and modify. To explore this, we experiment with a couple of multiplicative update schemes (detailed in App. B) across different layers of the model: the outer (unembedding) layer (Experiment 1), $Q$ matrices of atttention layers (Experiment 2) or all normalization layers (Experiment 3). We distinguish between low-rank (LoRa) - rank-$1$ in fact (Experiment 1), broadcast (Experiment 2) and full (Experiment 3) updates; LoRa multiplies the weight matrix by a low rank matrix, broadcast multiplies pointwise by a matrix with identical rows and full multiplies by an update matrix directly. For full details refer to App. B. We have deliberately used sparse (low rank, few layer) updates, optimizing over only a miniscule fraction of the parameters (between $\approx 4K$ and $\approx 266K$). Indeed, BBO poorly scales up, and the number of evaluations to reach some optimum is known to increase linearly with the dimension [82], making low-budget BBO out-of-reach, except when the actual dimensionality of the problem (dimension of the manifold of interest) is much lower – as is probably the case here.
[…]
@@
Experiment 1 (rank-1, output layer updates).
In a first series of experiments on Llama3.1-8B (resp. Llama3.1-8B-Instruct) we test BBoxER performance on GSM8k-test, mostly with 1 seed (sometimes 2). Figure 1 illustrates the results: while D-CMA was our favorite choice, we ran other BBO algorithms to validate the robustness of the BBoxER approach.
Experiment 2 (broadcast updates on attention layers).
@@
In this setting, we again train a range of BBO algorithms on the GSM8K training set on Llama3.1-8B (Base) using  [83], to explore out-of-distribution generalization capabilities of BBoxER with evaluations on SVAMP [84] and GSM+[55], in addition to GSM8k-test. We experiment with broadcast updates on $Q$-matrices of the 1st, 8th or all attention layers. We have added bet-and-run variants which enjoy particularly strong generalization guarantees (see  Sec. 4).
@@
Tab. 1 shows the results for different settings. Each line without std.dev. ($\pm$) corresponds to one single run, and all runs use a budget $b=150$ except otherwise specified; on the first attention layer, OneFifth was run with 3 seeds and D-CMA and CoLengler with 2 seeds each. Except for the “no few shot” line, default $8$-shot evaluation is used on GSM8k. The results demonstrate the stability of BBoxER in terms of learning capabilities, and a notable performance improvement of up to $7\%$ in distribution (GSM8k test) and $5\%$ in out-of-distribution transfer to GSM+. Additional results are reported in Tab. 5 in App. I.
@@
Experiment 3 (full update of normalization layers). In this experiment, D-CMA is used to directly optimize all normalization layers of Llama3.1-8B(-Instruct) and Qwen2.5-3B-Instruct (no low-rank and no broadcast involved) on GSM8k. As in Experiment 2, we test both in distribution (GSM8k-test), and on various other benchmark datasets (out of distribution). We report the results averaged over 5 different runs with standard deviation. The results (Tab. 2) indicate that running BBoxER improves in distribution performance and generally preserves it on very different tasks (e.g., Hellaswag) and sometimes even improves it (e.g., on ARC). BetAndRun results are reported in Tab. 6 and additional experiments using a larger training set in Tab. 7 (both in App. I).
@@
Llama3.1-8B-Instruct-BBoxER (b=150) 77.79$\pm$0.41 55.11$\pm$0.15 37.16$\pm$0.53 79.16$\pm$0.32 55.48$\pm$0.17 25.00$\pm$3.54 79.98$\pm$0.05
Llama3.1-8B-Instruct-BBoxER (b=300) 78.57$\pm$0.41 55.33$\pm$0.53 37.56$\pm$0.51 79.64$\pm$0.12 55.61$\pm$0.15 21.00$\pm$3.00 79.94$\pm$0.04
@@
Qwen-2.5-3B-instruct-BBoxER (b=150) 82.90$\pm$0.37 62.27$\pm$0.45 42.48$\pm$0.41 72.57$\pm$0.19 47.59$\pm$0.31 38.00$\pm$2.45 75.00$\pm$0.06
Qwen-2.5-3B-instruct-BBoxER (b=300) 83.55$\pm$0.36 61.94$\pm$0.64 41.32$\pm$0.95 72.62$\pm$0.17 47.76$\pm$0.22 36.50$\pm$3.00 75.02$\pm$0.09

### v2

@@ 4 Theoretical analysis
Analyzing BBoxER within a unified framework through its information bottleneck that compresses the dataset into a sequence of comparisons, allows us to state and prove strong and precise bounds on its ability to preserve generalization, its robustness to poisoning, its privacy guarantee, and its robustness against extraction attacks. For mathematical background see App.˜C, additional results in App.˜E, and proofs in App.˜G.
4.1 Overfitting bounds
In this section, we adress generalization by bounding the gap between empirical and true loss in terms of the variability of the algorithm’s internal state. Notably, these bounds are governed by the optimization path, with no dependence on the number of model parameters optimized during the retrofitting process.
Let $L(x)=L(modified(m_{0},x))$ denote the true loss over dataset distribution $\cal D$}, and $\widehat{L}(x)=\widehat{L}(modified(m_{0},x))$ the empirical loss on sampled dataset $D$. Let $N(\omega,a,b)$ denote the number of possible internal states over all possible datasets in $\cal D$}, then from ˜5 in Alg.˜1 we have:
@@ 4.1 Overfitting bounds
We fix the algorithm $a$ seed $\omega$ (the stochastic case is discussed in Sec.˜E.2) and study how the number of distinct possible internal states affects generalization. We defer the general case Thm˜1 to App.˜E, showing that the generalization gap is controlled by the number of reachable configurations in Equation˜4. It also provides a uniform bound over all intermediate models $(m_{1},\dots,m_{b})$, ensuring that empirical losses remain close to the true risk throughout optimization. The bound is obtained using mild assumptions, only requiring that the that empirical loss is a good proxy for the true loss, a standard concentration-bound hypothesis in classical learning theory [52, 53] (see Sec.˜C.1).
We present below the practical case where the individual losses are bounded in [0,1], as with accuracy or user preference signals $\in\{0,1\}$, and where data points are sampled independently from each other.
@@
(informal) Under standard concentration assumptions (Equation˜A.1), if individual losses are in $[0,1]$ and Hoeffding’s inequality applies (Eq. A.2) with dataset size s, then:
$P_{\omega_{D}}\left(|\widehat{L}(\widehat{x})-L(\widehat{x})|\geq\epsilon\right)\leq 2\cdot\left(\prod_{i=1}^{b}k_{i}\right)\cdot\exp(-2s\epsilon^{2})=2\cdot\exp\left(\ln 2\cdot\sum_{i=1}^{b}\log_{2}(k_{i})-2s\epsilon^{2}\right)$ (5)
Most BBO algorithms used within BBoxER bound the average $k_{i}$ by $2$ or $3$. Interpreting $\log_{2}(k_{i})$ as a budget in bits means that to avoid overfitting we can afford a number of iterations $b$ proportional to $s$ divided by the bits used per step. Precisely, if we assume $\forall i,k_{i}=2$, then for an error gap $\epsilon$ and confidence $\delta$, we have from Corollary˜1 the following formula (numerical values can be found in Sec.˜E.7):
$P_{\omega_{D}}\left(|\widehat{L}(\widehat{x})-L(\widehat{x})|\geq\epsilon\right)\leq\delta\Leftrightarrow b\leq\frac{\ln(\delta)+2s\epsilon^{2}}{\ln(2)}-1$ (6)
Bounds on additional BBO algorithms. Our analysis extends to a wide range of evolutionary BBO methods. In particular, Sec.˜E.4 shows that so called Bet-And-Run variants - running $N$ parallel runs each with a budget of $b/N$ - can reduce overfitting and scale the total budget while preserving generalization. Sec.˜E.5 and Sec.˜E.6 provide analogous bounds for Population-Based Evolution Strategies (e.g., CMA, DiagonalCMA), Differential Evolution, and Particle Swarm Optimization.
@@ 4.2 Privacy and poisoning
We now consider the two key risks associated with model post-training: (a) poisoning, where a small number of corrupted inputs can steer the outcome, and (b) privacy leakage, where individual training samples or - in the case of A/B testing - individual user preferences - might be inferred from the final model. Standard fine-tuning methods, including privacy-aware approaches such as differentially private stochastic gradient descent (DP-SGD), rely on batch-level gradient signals that inherently encode information on the underlying data. BBoxER  on the other hand, relies only on aggregate values over the entire dataset, namely on the outcomes of model comparisons $(choice_{1},\dots,choice_{b})$. Each $choice_{i}$ represents a compressed signal derived from the full dataset $D$. For a training dataset $D$, in this context we define preference signals to be the sample-level decisions that lead to aggregate decision $choice_{i}$ over $D$. In the case of A/B testing on prompt/output pairs, these would be the binary user preferences that aggregate to a model preference. In the case of more general metrics, the preference signal for a sample is defined as the result of models comparison over the subdataset {sample}.
@@
To analyze BBoxER, we consider a simplified setting, where $\forall i,k_{i}=2$, and in which two models $\{0,1\}$ are compared by majority vote over $s$ independent preference signals $r_{i}\in\{0,1\}$. The observed frequency is $f=\tfrac{1}{s}\sum_{i=1}^{s}r_{i}$. An adversary may corrupt up to $m$ values among the $r_{i}$, yielding a perturbed frequency $f^{\prime}$ such that $|f^{\prime}-f|\leq m/s$. This follows the standard poisoning model [54], where the adversary is computationally unbounded and fully informed, but constrained in the number of values it can alter. Lem.˜1 in App.˜G bounds the probability that such perturbations alter the comparison outcome.
@@
Using a union bound argument, we extend this analysis to BBoxER (Alg.˜1), which performs $b$ consecutive comparisons. In iteration $j\in\{1,\dots,b\}$, the algorithm collects binary preferences $r_{j,1},\dots,r_{j,s}$, aggregates them into $f_{j}=\tfrac{1}{s}\sum_{i=1}^{s}r_{j,i}$, and selects a model based on whether $f_{j}>1/2$. Under poisoning, the adversary may shift $f_{j}$ to $f^{\prime}_{j}$ with $|f^{\prime}_{j}-f_{j}|\leq m/s$, independently at each round. The following theorem bounds the impact of these perturbations on the final output.
Theorem 2 (Robustness of BBoxER under poisoning).
[…]
$P\!\left(output_{1}\neq output_{2}\right)\;\leq\;b\frac{2m+1}{\sqrt{2s\pi}}.$ (7)
Differential privacy guarantee Let $D_{1}$ and $D_{2}$ be two datasets that differ by a single sample. We examine the effect of this difference on the FinalModel given by BBoxER, and we distinguish two cases:
$\bullet$ privacy of users prompts and outputs (like in A/B testing) We assume that the preference signals are provided directly by users, and that the datasets $D_{1}$ and $D_{2}$ can differ on the prompts or outputs. Since BBoxER relies solely on users preference signals to perform model comparisons, and does not access prompts or outputs, privacy is ensured by design and we have $(\varepsilon,\delta)=(0,0)$.
$\bullet$ privacy of the entire sample We assume that the $D_{1}$ and $D_{2}$ can differ on an entire sample, which impacts its prefence signal. This is the case when the preference signal needs to be computed from the sample before aggregating and passing it to the algorithm. In this scenario, we obtain $\varepsilon=0$, while $\delta\;\propto\;\frac{b}{\sqrt{s}}$, following the poisoning bound from Thm˜2 with $m=1$.
Thm˜2 yields two immediate consequences: First, unless $m=\Omega(\sqrt{s})$, the retrofitting framework is resilient to poisoning attacks, and second, privacy is guaranteed by design, covering prompts, outputs, and user satisfaction signals without additional mechanisms. These points are further discussed in App.˜F. It is important to note that, while other privacy preserving methods such as DP-SGD [46] enforce privacy by perturbing gradients and consequently suffer from utility degradation as $\varepsilon$ becomes small, BBoxER’s compression-based guarantees remain robust. These structural differences highlight a fundamental advantage of our approach.
4.3 Robustness to extraction attacks
@@
Formally, let $m_{0}$ be fine-tuned on unrelated datasets $D_{1},\dots,D_{n}$. For text, "unrelated" can mean differing by one word out of five while remaining coherent. A dataset $D$ with $B$ bits may yield $n\approx 2^{B/k}$ such variants by altering one bit per $k$. We define vulnerability to data extraction as follows:
@@
This criterion can be tested practically: generate $n$ unrelated variants $D_{1},\dots,D_{n}$, fine-tune models on each $D_{i}$, and check whether each model assigns the highest likelihood to its own training set. If so, it indicates vulnerability to data extraction. We argue this is unlikely for retrofitting, due to limited capacity to encode training data. Thm˜5 in Sec.˜G.2 formalizes this, leading to the corollary below:
@@
For given $\omega,a,b$ s.t. $k_{i}\leq 2$ for all $i$, BBoxER is not vulnerable to extraction for ${\cal D}=(D_{1},\dots,D_{n})$ if $n>2^{b}$. More generally, BBoxER is not vulnerable for $\cal D$} if $n>\prod_{i}k_{i}$.
Thus, if the number of alternative datasets exceeds that of distinguishable outputs, no extractor can identify the training data. Since $2^{b}$ grows only exponentially with budget $b$, while the number of unrelated textual variants grows exponentially with dataset size (e.g., >400>400B bits for Wikipedia articles, >20>20M for War and Peace), extraction is very unlikely in practice.
@@ 5 Experimental results: Learning with Privacy and Robustness
The theoretical bounds in Sec.˜4 establish that comparison based BBOs in the framework of BBoxER offer strong generalization guarantees and inherent privacy. However, theoretical guarantees alone do not ensure that these methods are practically effective—only empirical evidence can reveal whether they actually lead to meaningful learning. To validate the capacity of BBoxER as an LLM retrofitting method, we focus on verifiable domains, like math, where performance improvements are both measurable and interpretable. We experiment with budgets $b$ determined theoretically for guaranteed generalization on the test set and we empirically examine how performance changes as the budget exceeds these thresholds.
@@ 5.1 Experiments
Experimental setup. We deploy a family of BBO algorithms with finite branching factors, described in App.˜D, using the default parameters in Nevergrad. We conduct experiments on both base and instruct variants. Retrofitting the base models allows us to verify that learning indeed occurs, while targeting the instruct models demonstrates that our method can yield further improvements even on top of models already fine-tuned for reasoning tasks. All experimental details can be found in App.˜H.
Datasets. Most of our experiments deploy BBoxER on GSM8K [56], a dataset composed of short grade-school math problems, consisting of $7,473$ problems for training and $1,319$ for testing. We train on the train split and test on GSM8k-test (in distribution) with standard 8-shot methodology (unless otherwise stated), in addition to various other benchmark datasets (out of distribution): MATH500 (4-shots) [57], Hellaswag (0-shots) [58], GSM+ test_mini (8-shots) [59], ARC Easy (0-shots) and ARC Challenge (0-shots) [60], AMC23 (0-shots). All benchmarks are evaluated with exact match using greedy decoding. The theoretical bounds in Sec.˜4 allow us to precisely determine the budget $b$ that guarantees in-domain generalization (as a function of dataset size): for GSM8k-train the permissible budget is between 100100-$150$ steps and approximately $3$ times larger for Bet-and-run variants like 3-OneFifth.
Choices for parameter updates. A key design choice in our framework is choosing which parameters the BBO algorithm should modify. To explore this, we experiment with a couple of multiplicative update schemes, across different layers of the model: the outer (unembedding) layer (Experiment 1), $Q$ matrices of attention layers (Experiment 2) or all normalization layers (Experiment 3). We distinguish between low-rank (LoRa) - rank-$1$ in fact (Experiment 1), broadcast (Experiment 2) and full (Experiment 3) updates; LoRa multiplies the weight matrix by a low rank matrix, broadcast multiplies pointwise by a matrix with identical rows and full multiplies by an update matrix directly. For full details, refer to App.˜B. We optimize a small number of parameters, between $\approx 4K$ and $\approx 266K$, as BBO algorithms are known to scale poorly unless the problem’s effective dimensionality is much lower [61], which is likely the case here.
Experiment 1 (rank-1, output layer updates). In a first series of experiments on Llama3.1-8B (resp. Llama3.1-8B-Instruct) we test BBoxER performance on GSM8k-test, mostly with 1 seed (sometimes 2). Figure˜1 illustrates the results: while D-CMA was our favorite choice, we ran other BBO algorithms to validate the robustness of the BBoxER approach.
Table 1: Experiment 2. Performance difference (after-before) BBoxER with Llama3.1-8B (Base) on GSM8k-train, evaluated on GSM8k-test (ID), (OOD) SVAMP and GSM+. Red: detrimental runs, $1\%$. Green: beneficial runs, $1\%$. Blue rows are bet-and-run cases.
@@
Experiment 2 (broadcast updates on attention layers). In this setting, we again train a range of BBO algorithms on the GSM8K training set on Llama3.1-8B (Base) to explore out-of-distribution generalization capabilities of BBoxER with evaluations on SVAMP [62] and GSM+[59], in addition to GSM8k-test. We experiment with broadcast updates on $Q$-matrices of the 1st, 8th, or all attention layers. We have added bet-and-run variants which enjoy particularly strong generalization guarantees (see Sec.˜4).
@@
Tab.˜1 shows the results for different settings. Each line without std.dev. ($\pm$) corresponds to one single run, and all runs use a budget $b=150$ except otherwise specified; on the first attention layer, OneFifth was run with 3 seeds and D-CMA and CoLengler with 2 seeds each. Except for the “no few-shot” line, the default $8$-shot evaluation is used on GSM8k. The results demonstrate the stability of BBoxER in terms of learning capabilities, and a notable performance improvement of up to $7\%$ in distribution (GSM8k test) and $5\%$ in out-of-distribution transfer to GSM+. Additional results are reported in Tab.˜6 in App.˜I.
@@
Experiment 3 (full update of normalization layers). In this experiment, D-CMA is used to directly optimize all normalization layers of Llama3.1-8B(-Instruct) and Qwen2.5-3B-Instruct (no low-rank and no broadcast involved) on GSM8k. As in Experiment 2, we test both in distribution (GSM8k-test), and on various other benchmark datasets (out of distribution). We report the results averaged over 5 different runs with the standard deviation. The results (Tab.˜3) indicate that running BBoxER improves in distribution performance and generally preserves it on very different tasks (e.g., Hellaswag). Additional experiments are described in App.˜I: BetAndRun results are reported in Tab.˜7, and experiments using a larger training set can be found in Tab.˜8 in App.˜I.
Table 2: Experiment 4 - BBoxER is robust to MIA. Comparison of BBoxER, SFT and DP-SFT in the setting of Experiment 3 with Llama3.1-8B-Instruct on GSM8K. Runs with comparable test accuracy are highlighted in the same color to compare robustness at similar accuracy. We see that both finetuning and DP-finetuning approaches lead to NLL metrics at least an order of magnitude larger than comparable BBoxER runs.
Method Diff-Full Diff-CoT-a Acc
ft-5epochs 2.25e-01 5.06e-01 +7.88
ft-2epochs 2.53e-01 3.99e-01 +3.79
ft-norm-5epochs 6.00e-03 4.62e-02 +0.45
DP-AdamW-eps=8 8.66e-01 2.17e-01 +2.05
BBoxER-norm-b=150 7.19e-04 2.90e-03 +0.15
BBoxER-norm-b=300 1.01e-03 3.68e-03 +1.97
BBoxER-norm-b=1200 6.33e-03 1.17e-02 +3.87
Experiment 4 (Robustness to data extraction). To assess the robustness of BBoxER against membership inference attacks (MIAs), we begin from the observation that most MIAs rely on the negative log-likelihood (NLL) of training samples as their primary signal to infer membership [63, 64, 65] - see also App.˜H for an exposition. Accordingly, we propose to measure the average of absolute differences in negative log-likelihood (NLL) between the base and the updated model for each training sample. We consider GSM8K as our training dataset, Llama-3.1-8B as our base model, and compute three metrics: Diff-full (NLL computed over the entire prompt), Diff-CoT-a (NLL computed over only the CoT and the answer), and the relative average exact match improvement on GSM8K-test from the base model. For a fair comparison, we evaluate BBoxER alongside standard fine-tuning and DP-AdamW, training all methods on GSM8K under different settings-varying budget, epochs, and parameter subsets (full details in App.˜H)—so that models are compared at similar utility. Our results, reported in Tab.˜2, demonstrate that at matched test accuracy, the NLL shifts under BBoxER are orders of magnitude smaller than under fine-tuning or DP-AdamW. This provides empirical confirmation of our theoretical claim in Sec.˜4.3: BBoxER is significantly more robust to extraction and membership inference attacks.
@@
[…]
@@
[…]